Let a be the number such that the area to the right of z = a is 0.21. Find the area between z = –a and z = a.

A. 0.21 B. 0.42
C. 0.58 D. None of the above.
Answer #1

Given,

P(Z > z) = 0.21

That is

P(Z < z) = 1 - 0.21

P(Z < z) = 0.79

Now,

P(-z < Z < z) = p(Z < z) - P(Z < -z)

= P(Z < z) - ( 1 - P(Z < z) )

= 0.79 - ( 1 - 0.79)

= 0.58